All discount rates are 6% per period unless otherwise stated. All cash flows are end-of-period. Answers are in parens, from Excel, so they may differ slightly from table answers due to rounding.

  • Consider CF’s of $1500, 1400, 1300, etc, down to CF $600. (10 CF’s total.) What’s PV? ($8,079.90)

  • Would you prefer $500 per period for 10 periods, or CF’s of $100,$200, $300...to $1000 in the 10th period? ($502.20)
Answer #1

a. NPW of cash flow = 1500 * (P/A,6%,10) - 100 * (P/G,6%,10)

= 1500 *7.360087 - 100 *29.602321

= 8079.90

b.

AW of cash flow for option 1 = 500

AW of cash flow for option 2 = 100 + 100 * (A/G,6%,10)

= 100 + 100 * 4.022007

= 502.20

AW of cash flow for option 2, so it should be selected